# 11footnotetext: Corresponding authors22footnotetext: Project lead

### v1

@@ Abstract
Fine-tuning large-scale text-to-image diffusion models for various downstream tasks has yielded impressive results. However, the heavy computational burdens of tuning large models prevent personal customization. Recent advances have attempted to employ parameter-efficient fine-tuning (PEFT) techniques to adapt the floating-point (FP) or quantized pre-trained weights. Nonetheless, the adaptation parameters in existing works are still restricted to FP arithmetic, hindering hardware-friendly acceleration. In this work, we propose IntLoRA, to further push the efficiency limits by using integer type (INT) low-rank parameters to adapt the quantized diffusion models. By working in the integer arithmetic, our IntLoRA offers three key advantages: (i) for fine-tuning, the pre-trained weights are quantized, reducing memory usage; (ii) for storage, both pre-trained and low-rank weights are in INT which consumes less disk space; (iii) for inference, IntLoRA weights can be naturally merged into quantized pre-trained weights through efficient integer multiplication or bit-shifting, eliminating additional post-training quantization. Extensive experiments demonstrate that IntLoRA can achieve performance on par with or even superior to the vanilla LoRA, accompanied by significant efficiency improvements. Code is available at https://github.com/csguoh/IntLoRA.
@@ 1 Introduction
Recently, large-scale text-to-image (T2I) diffusion models (Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022; Podell et al., 2023) have shown promising capabilities for image generation. Taking advantage of the strong generative prior of pre-trained parameters, a range of downstream adaptation applications have emerged, including subject-driven generation (Ruiz et al., 2023), style-customized generation (Sohn et al., 2023), and controllable generation (Zhang et al., 2023). However, fully fine-tuning these large models for personalized customization poses challenges on consumer-level GPUs, as well as the costs of storing the weight for each downstream task.
To facilitate efficient adaptation, recent advances (Qiu et al., 2023; Liu et al., 2023b) have introduced parameter efficient fine-tuning (PEFT) (Hu et al., 2021; Houlsby et al., 2019; Jia et al., 2022) that focuses on fine-tuning a limited number of parameters for downstream tasks. Despite the reduction in trainable parameters, current PEFT methods predominantly rely on floating-point (FP) arithmetic, which can be inefficient for practical applications. For instance, using large FP pre-trained weights during fine-tuning can result in significant GPU memory overhead. Additionally, both the pre-trained weights and the adaptation parameters remain in FP format, consuming substantial disk space. Although some PEFT methods propose weight merging techniques to mitigate the extra inference cost associated with the adaptation parameters, they still fall short in obtaining reduced latency beyond the pre-trained large models.
On the other hand, neural network quantization (Nagel et al., 2021; 2020; Esser et al., 2019), which can transform the trained FP parameters into low-bit integer (INT) representations, is a prominent technique for accelerating deep learning models. Thus, integrating PEFT with quantization techniques holds promise for enhancing the efficiency of downstream adaptations. To this end, some pioneer works (Dettmers et al., 2024; Qin et al., 2024) attempt downstream adaptation by directly fine-tuning the quantized INT pre-trained weights with FP PEFT parameters. However, it remains an open challenge for quantization-aware adaptation. Specifically, since the adaptation weights are still in FP, it is inevitable to convert the quantized pre-trained weights back to FP for arithmetic consistency for subsequent weight merge. As a result, it necessitates additional post-training quantization (PTQ) to quantize the merged weights for practical employment, which is pipeline-complicated and incurs a performance drop when the bit-width is low (see Fig. 1).
To some extent, the primary technical bottleneck in bridging PEFT and quantization arises from the arithmetic inconsistency between the quantized pre-trained weights and the adaptation weights. Specifically, the FP adaptation weights necessities the conversion of the quantized pre-training weights from INT to FP for the subsequent weight fusion. To address this inconsistency, a potential solution is to also transfer the adaptation weights to integer arithmetic. In this way, all weights during fine-tuning are quantized, allowing to store only INT weights, as well as convenient weight merging during inference. Despite these promising properties, it is non-trivial to accurately quantize the adaptation weights for satisfactory performance. For example, while zero initializing low-rank weights are advantageous for fine-tuning (Hu et al., 2021), it poses quantization challenges due to substantial quantization errors from small values. Furthermore, the additive form of the original LoRA forces the pre-trained and adaptation weights to share the same quantizer for subsequent weight merging, which restricts the available parameter space for adaptation.
@@
Figure 1: Left: (a) The arithmetic inconsistency between the pre-trained and adaptation weights results in the merged weights still in FP. Consequently, additional PTQ is needed for low-bit inference. (b) Our approach allows PEFT to operate directly on INT arithmetic, ensuring the merged weights seamlessly in INT format and streamlining the whole process. Right: The utilization of PTQ on the downstream adapted weights leads to severe performance degradation under low bit-width quantization, which makes the tuning-then-quantization paradigm less general.
@@
To address the above challenges, we propose IntLoRA, which employs INT low-rank parameters to adapt the quantized diffusion models. In detail, we introduce the Adaptation-Quantization Separation (AQS) technique, which employs a task-agnostic auxiliary matrix to enable quantization-friendly low-rank parameters without disrupting the gradient trajectory of the original LoRA. Additionally, we present the Multiplicative Low-rank Adaptation (MLA), which reformulates the mathematical structure of LoRA from addition to multiplication. This remains mathematically equivalent to the original but allows for independent optimization of adaptation weights, eliminating the need to share the same quantizer as the pre-trained weights. Furthermore, we develop the Variance Matching Control (VMC) mechanism that aligns variances of the pre-trained and auxiliary matrices, controlling the adaptation distribution for more efficient log2-quantization. For implementation, we provide two versions of IntLoRA, i.e., IntLoRAMUL, and IntLoRASHIFT. The IntLoRAMUL learns quantized low-rank parameters and can be seamlessly merged with quantized pre-trained weight through integer multiplication, while IntLoRASHIFT introduces log2-quantization and operates by bit-shifting the quantized pre-trained weights for downstream adaptation. We evaluate our IntLoRA on prevalent T2I adaptation tasks, including subject-driven generation (Ruiz et al., 2023), style-customized generation (Sohn et al., 2023) and controllable generation (Zhang et al., 2023). Extensive experiments demonstrate that IntLoRA represents a novel diffusion fine-tuning paradigm with impressive efficiency and favorable performance.
The contribution of this work can be summarized as follows: (i) we introduce IntLoRA, which achieves integer PEFT to address the arithmetic inconsistency, thereby advancing the efficiency of diffusion model adaptations; (ii) we propose the adaptation-quantization separation to facilitate quantization-friendly pre-trained weights, and further develop the multiplicative low-rank adaptation for independent quantizers, complemented by variance matching control for effective distribution manipulation; (iii) our IntLoRA enables the adaptation of quantized diffusion models through hardware-friendly integer multiplication or bit-shifting, resulting in significant efficiency gains in fine-tuning, storage, and inference. Extensive experiments validate the superiority of our method.
@@ 2 Related Work
Parameter-efficient fine-tuning of diffusion models.
@@
In order to reduce the fine-tuning cost of large models, parameter-efficient fine-tuning (PEFT) has recently gained great interests (Lian et al., 2022; Chavan et al., 2023; Li & Liang, 2021; He et al., 2021; Jie & Deng, 2023). For example, some pioneering work (Zaken et al., 2021) proposes to fine-tune only a fraction of the pre-trained weights. To achieve better performance, prompt-based methods (Jia et al., 2022) append learnable prompts to modify the input space. Adapter-based methods (Houlsby et al., 2019; Chen et al., 2022) employ additional bottleneck structures as bypass branches for adaptation. LoRA (Hu et al., 2021), which adopts low-rank matrices to learn weight updates for downstream tasks, has become a popular pipeline for diffusion model adaptation. Moreover, OFT (Qiu et al., 2023) uses orthogonal constraints to preserve the pairwise angles between neuron vectors to maintain the hypersphere energy. COFT (Liu et al., 2023b) further uses butterfly factorization to formulate denser orthogonal matrices using fewer parameters. In this work, we mainly focus on LoRA since it has been widely applied and can be merged into pre-trained weights without increasing the inference cost.
Network quantization of diffusion models.
@@
Quantization (Nagel et al., 2021) is an effective technique to reduce model storage and inference costs, and can be categorized into quantization-aware training (QAT) (Jacob et al., 2018; Li et al., 2024; 2022; Xu et al., 2023a) and post-training quantization (PTQ) (Wang et al., 2023a; Nahshan et al., 2021; Li et al., 2021; Wei et al., 2022; Liu et al., 2023a; Huang et al., 2024a). In the context of diffusion model quantization, existing works mainly focus on PTQ because of the significant overhead of retraining diffusion models. For example, PTQ4DM (Shang et al., 2023) makes the first attempt to quantize diffusion models to 8 bits. After that, Q-Diffusion (Li et al., 2023) further achieves improved performance and lower bit-width. In addition, PTQD (He et al., 2024) eliminates quantization noise through correlated and residual noise correction. EfficientDM (He et al., 2023) introduces LoRA to fine-tune the pre-trained model to allow comparable performance with QAT. TFMQ-DM (Huang et al., 2024b) observes the impact of quantization errors on the time-step representation and proposes to quantize the time-embedding layer individually for better performance. However, the network quantization only enables efficient inference at low bit-widths on the original task, failing to handle various downstream tasks.
Joint adaptation and quantization.
@@
The work aims to achieve both parameter-efficient adaptations as well as storage and inference-efficient quantization, to further push the efficiency limits of diffusion model fine-tuning. However, this also poses additional challenges, such as back-propagating gradients in quantized values and optimizing the learnable parameters with quantization restrains. Existing related works have explored mainly LLMs, but it is still far from a satisfactory solution. Specifically, QLoRA (Dettmers et al., 2024) proposes to quantize the LLMs before fine-tuning the LLMs with LoRA. Despite the reduced GPU usage during training due to the import of only the quantized model, QLoRA does not maintain quantized at inference since the quantized weights need to be converted to FP again so that to add the LoRA weights. To solve this problem, QA-LoRA (Xu et al., 2023b) develops a group-wise quantization through sharing parameters across channels, but at the cost of impairing the adaptation ability as well as the need to store FP LoRA weights. IR-QLoRA (Qin et al., 2024) analyzes the entropy loss of quantization from an information theory view, but it also needs to convert the quantized weights back to FP during inference.
@@ 3 Preliminary
The main idea of low-rank adaptation (LoRA)  (Hu et al., 2021) comes from the fact that the learned incremental matrix during downstream adaptation usually possesses low-rank properties. To this end, LoRA introduces a low-rank matrix $\Delta\mathbf{W}$ to learn the weight increments for adapting the pre-trained weights $\mathbf{W}\in\mathbb{R}^{C_{out}\times C_{in}}$ to downstream tasks. In implementation, the $\Delta\mathbf{W}$ is formulated as the matrix multiplication of two low-rank matrices $\mathbf{A}\in\mathbb{R}^{C_{out}\times d}$ and $\mathbf{B}\in\mathbb{R}^{d\times C_{out}}$, where the inner dimension $d$ is the pre-defined rank. During fine-tuning, the pre-trained weight $\mathbf{W}$ is frozen and only $\mathbf{A},\mathbf{B}$ is trainable. Since $d\ll\min\{C_{in},C_{out}\}$, the number of trainable parameters can be very small compared to full fine-tuning. The output during the downstream fine-tuning is calculated as $\mathbf{y}=\mathbf{W}\mathbf{x}+\lambda\cdot(\mathbf{A}\mathbf{B})\mathbf{x}$, where $\lambda$ is the LoRA scale to adjust the control strength. During inference, the task-specific $\mathbf{AB}$ can be naturally merged into the pre-trained weights, i.e., $\mathbf{W^{\prime}}=\mathbf{W}+\lambda\cdot\mathbf{A}\mathbf{B}$, without increasing additional computational cost.
Even though the PEFT techniques can alleviate training costs, they still struggle to reduce the inference latency. To allow for accelerated inference, network quantization is a common practice that converts the FP weights to hardware-efficient low-bit INT integers. Formally, given a FP tensor $\mathbf{X}$, the target bit-width $b$, the quantization process can be defined as:
@@ 4 Methodology
4.1 Bridging Efficient-adaptation and Quantization
Although parameter-efficient adaptation and network quantization each contribute to improved fine-tuning and inference efficiency, it remains an unexplored challenge to effectively combine them. Existing work such as QLoRA (Dettmers et al., 2024) employs the low-rank $\mathbf{AB}$ to fine-tune quantized weights $\hat{\mathbf{W}}$ to reduce the memory footprint of loading a full-precision model. After training, merging the low-rank FP parameters and quantized INT weights results in the adapted weights $\mathbf{W^{\prime}}$ which is reverted to FP again. Consequently, additional PTQ on $\mathbf{W^{\prime}}$ is needed for inference efficiency. When the bit width is low, e.g., 4-bit, this PTQ can significantly degrade performance.
In this work, we attribute the main difficulty in effectively marrying PEFT and quantization to arithmetic inconsistency. Specifically, the FP $\mathbf{AB}$ forces the quantized INT $\hat{\mathbf{W}}$ to revert to FP during weight merging. To bridge the arithmetic gap, a feasible solution is to convert the low-rank parameters to integer form $\hat{\mathbf{AB}}$. In this way, the same typed $\hat{\mathbf{W}}$ and $\hat{\mathbf{AB}}$ can be seamlessly merged during inference without additional PTQ. However, several technical challenges arise when performing PEFT on integer arithmetic. First, the $\mathbf{AB}$ in the original LoRA is zero-initialized to preserve the knowledge of the pre-trained weights. Although helpful for fine-tuning, this initialization complicates the quantization process. For instance, the all-zero distribution requires a separately designed quantizer at the beginning of tuning. In addition, quantizing small values also produces large relative errors, leading to inaccurate gradient optimization. Second, the vanilla LoRA merges the FP $\mathbf{AB}$ and $\mathbf{W}$ using addition. When both $\mathbf{AB}$ and $\mathbf{W}$ are quantized, it is essential to ensure that they share identical quantization parameters to enable PTQ-free weight merging. This constraint restricts the feasible parameter space, thus limiting the adaptation ability.
Figure 2: The overall adaptation pipeline of our proposed IntLoRA.

### v3

# IntLoRA: Integral Low-rank Adaptation of Quantized Diffusion Models

@@ Abstract
Fine-tuning pre-trained diffusion models under limited budgets has gained great success. In particular, the recent advances that directly fine-tune the quantized weights using Low-rank Adaptation (LoRA) further reduces training costs. Despite these progress, we point out that existing adaptation recipes are not inference-efficient. Specifically, additional post-training quantization (PTQ) on tuned weights is needed during deployment, which results in noticeable performance drop when the bit-width is low. Based on this observation, we introduce IntLoRA, which adapts quantized diffusion models with integer-type low-rank parameters, to include inference efficiency during tuning. Specifically, IntLoRA enables pre-trained weights to remain quantized during training, facilitating fine-tuning on consumer-level GPUs. During inference, IntLoRA weights can be seamlessly merged into pre-trained weights to directly obtain quantized downstream weights without PTQ. Extensive experiments show our IntLoRA achieves significant speedup on both training and inference without losing performance. Code is available at https://github.com/csguoh/IntLoRA.
Low-rank Adaptation, Network Quantization, Diffusion Models
@@ 1 Introduction
Recently, large-scale text-to-image diffusion models (Rombach et al., 2022; Saharia et al., 2022; Podell et al., 2023) have shown promising capabilities for image generation. Taking advantage of the strong generative prior of pre-trained parameters, a range of downstream adaptation applications have emerged, such as subject-driven generation (Ruiz et al., 2023), style-customized generation (Sohn et al., 2023), and controllable generation (Zhang et al., 2023). However, fully fine-tuning large pre-trained models for downstream tasks poses challenges on consumer-level GPUs. For instance, only loading the FP32 FLUX.1-dev (BlackForestLabs, 2024) weights into GPUs can consume over 24GB of memory, let alone subsequent fine-tuning. Therefore, the huge fine-tuning costs hinder personalized diffusion model customization.
@@
Figure 1: (a) The arithmetic inconsistency between the pre-trained and adaptation weights leads to the merged weights still in FP16. Consequently, additional PTQ is needed for low-bit inference. (b) Our IntLoRA allows to work directly on INT4 arithmetic, ensuring the merged weights seamlessly in INT4 format and streamlining the whole process.
To facilitate efficient training, recent advances have introduced parameter efficient fine-tuning (PEFT) (Houlsby et al., 2019; Jia et al., 2022) techniques, such as LoRA (Hu et al., 2021), to fine-tune a limited number of parameters. With the reduced gradient and optimizer states, they can achieve comparable or even better adaptation performance than fully fine-tuning. More recently, some works (Dettmers et al., 2024; Qin et al., 2024) have successfully married PEFT and network quantization to allow the low-rank adaptation directly on the quantized weights (as shown in Fig. 1(a)). Through reducing the bit-widths, the GPU costs during fine-tuning are further decreased.
Although the reduced training costs have facilitated user customization, obtaining an inference-aware tuning recipe remains an open challenge. Specifically, existing methods predominantly employ floating-point (e.g., FP16) low-rank parameters during training, as a result, it is inevitable to convert the quantized pre-trained weights back to FP16 for arithmetic consistency to merge low-rank weights into pre-trained weights. During test-time, this pipeline necessitates additional post-training quantization (PTQ) on the FP16 merged weights for accelerated inference, which is pipeline-complicated and incurs significant performance drop when the bit-width is low (see Fig. 2).
Figure 2: The utilization of PTQ on the downstream merged weights leads to severe performance degradation under low bit-width quantization.
To address these challenges, a potential solution is to also transfer the adaptation weights to integer arithmetic. In this way, all weights during fine-tuning are in integers, thus ensuring the merged weights naturally being quantized. Despite these promising properties, it is non-trivial to accurately quantize the low-rank weights. For example, while zero initializing low-rank weights are advantageous for fine-tuning (Hu et al., 2021), it poses quantization challenges due to substantial quantization errors from small values. Furthermore, the additive form of the original LoRA forces the pre-trained and adaptation weights to share the same quantizer for seamless weight merging, which restricts available parameter space during fine-tuning.
@@
In this work, we propose IntLoRA, which achieves integral low-rank parameters for both training and inference efficient diffusion models. In detail, we introduce the Adaptation-Quantization Separation (AQS) technique, which employs a task-agnostic auxiliary matrix to enable quantization-friendly low-rank parameters without disrupting the gradient trajectory of the original LoRA. Additionally, we present the Multiplicative Low-rank Adaptation (MLA), which reformulates the mathematical structure of LoRA from addition to multiplication. This remains mathematically equivalent to the original but allows for independent optimization of adaptation weights. Furthermore, we develop the Variance Matching Control (VMC) to align the pre-trained and auxiliary matrices. For implementation, we provide two versions, i.e., IntLoRAMUL, and IntLoRASHIFT. The IntLoRAMUL learns quantized low-rank parameters and can be seamlessly merged through integer multiplication, while IntLoRASHIFT introduces log2-quantization and operates by bit-shifting the quantized weights for downstream adaptation. We evaluate our IntLoRA on various diffusion personalization tasks. Extensive experiments show that IntLoRA presents impressive efficiency and performance.
@@ 2 Related Work
Parameter-efficient fine-tuning of diffusion models. In order to reduce the fine-tuning cost of large models, parameter-efficient fine-tuning (PEFT) has recently gained great interests (Lian et al., 2022; Chavan et al., 2023; Li & Liang, 2021; He et al., 2021; Jie & Deng, 2023). For example, prompt-based methods (Jia et al., 2022) append learnable prompts to modify the input space. Adapter-based methods (Houlsby et al., 2019; Chen et al., 2022) employ additional bottleneck structures as bypass branches for adaptation. Moreover, LoRA (Hu et al., 2021) adopts low-rank matrices to learn the weight updates for downstream tasks. In this work, we mainly focus on LoRA since it has been widely applied in diffusion model fine-tuning and can be merged into pre-trained weights without increasing inference costs.
@@
Network quantization of diffusion models. Quantization (Nagel et al., 2021) is an effective technique to speed deep-learning models and can be categorized into quantization-aware training (QAT) (Jacob et al., 2018; Li et al., 2024, 2022; Xu et al., 2023a) and post-training quantization (PTQ) (Wang et al., 2023; Nahshan et al., 2021; Li et al., 2021; Wei et al., 2022; Liu et al., 2023; Huang et al., 2024a). In the context of diffusion model quantization, existing works mainly focus on PTQ because of the significant overhead of retraining diffusion models. For example, PTQ4DM (Shang et al., 2023) makes the first attempt to quantize diffusion models to 8 bits. After that, Q-Diffusion (Li et al., 2023) further achieves improved performance and lower bit-width. EfficientDM (He et al., 2023) introduces LoRA to fine-tune the pre-trained model to allow comparable performance with QAT. TFMQ-DM (Huang et al., 2024b) proposes to quantize the time-embedding layer individually for better performance.
@@
Joint PEFT and quantization for efficient fine-tuning. Benefiting from the scaling law, the pre-trained models have become increasingly large, which makes even loading models challenging. To allow fine-tuning on consumer-level GPUs for user customization, some work attempts to apply PEFT techniques directly on the quantized pre-trained weights. Specifically, QLoRA (Dettmers et al., 2024) proposes to quantize the LLMs before fine-tuning the LLMs with LoRA. Despite the reduced GPU usage during training due to the import of only the quantized model, QLoRA does not maintain quantized at inference since the quantized weights need to be converted to FP16 again so as to be merged with the LoRA weights. QA-LoRA (Xu et al., 2023b) develops a group-wise quantization through sharing parameters across channels but at the cost of impairing the adaptation ability. IR-QLoRA (Qin et al., 2024) analyzes the entropy loss of quantization from an information theory view, but it also needs to convert the quantized weights back to FP16 during inference.
@@ 3 Preliminary
The LoRA (Hu et al., 2021) introduces a low-rank matrix $\Delta\mathbf{W}$ to learn the weight increments for adapting the pre-trained weights $\mathbf{W}\in\mathbb{R}^{C_{out}\times C_{in}}$ to downstream tasks. In implementation, the $\Delta\mathbf{W}$ is formulated as the matrix multiplication of two low-rank matrices $\mathbf{A}\in\mathbb{R}^{C_{out}\times d}$ and $\mathbf{B}\in\mathbb{R}^{d\times C_{in}}$, where the inner dimension $d$ is the pre-defined rank. During fine-tuning, the pre-trained weight $\mathbf{W}$ is frozen and only the low-rank $\mathbf{A},\mathbf{B}$ are trainable. Since $d\ll\min\{C_{in},C_{out}\}$, the number of trainable parameters can be very small compared to full fine-tuning, thus reducing the GPU footprint of gradients and optimizer states. The output during downstream fine-tuning is calculated as $\mathbf{y}=\mathbf{W}\mathbf{x}+\lambda\cdot(\mathbf{A}\mathbf{B})\mathbf{x}$, where $\lambda$ is the LoRA scale to adjust the control strength. During inference, the task-specific $\mathbf{AB}$ can naturally be merged into the pre-trained weights, i.e., $\mathbf{W^{\prime}}=\mathbf{W}+\lambda\cdot\mathbf{A}\mathbf{B}$, without increasing additional costs.
Even though the LoRA can alleviate training costs through reduced gradients and optimizer states, it still needs to load huge FP16 pre-trained weights. Given the increasing pre-trained model size, it becomes impractical to only use LoRA to fine-tune the diffusion models on consumer-level GPUs. To further reduce the training memory, recent advancements (Dettmers et al., 2024; Xu et al., 2023b; Qin et al., 2024) have introduced network quantization to allow direct fine-tuning on the integer weights. Formally, given a tensor $\mathbf{X}$, the target bit-width $b$, the quantization process can be defined as:
@@
Despite current methods allow user to train customized models under a low memory budget, they all require additional PTQ on the fine-tuned weights for fast inference, which leads to noticeable performance degradation when the quantization bit-width is low.
Figure 3: Before tuning, we propose the Adaptation Quantization Separation (AQS) to incorporate auxiliary matrix into pre-trained weights and low-rank weights for zero-initialized but quantization-friendly distribution. Then, the Multiplicative Low-rank Adaptation (MLA) is used to reformulate additive LoRA into the product of the “pre-training term” and the “adaptation term”. At last, we introduce the Variance Matching Control (VMC) to adjust the distribution of the adaptation term by modulating the auxiliary matrix. After tuning, we use hardware-friendly integer multiplication or bit shifting to directly generate quantized merged weights without additional PTQ. The detailed algorithm is given in Appendix A.
@@ 4 Methodology
In this work, we aim to remove the additional PTQ of the merged weights by introducing integer-type low-rank parameters during fine-tuning. In this way, both the $\mathbf{AB}$ and $\mathbf{W}$ are in the same arithmetic type, thus ensuring the merged $\mathbf{W}^{\prime}$ is naturally already quantized. However, several technical challenges arise when transferring LoRA to integer arithmetic. First, the $\mathbf{AB}$ in the original LoRA is zero-initialized to ensure the behavior of the model is similar to the pre-trained one at the beginning of training. Although helpful for fine-tuning, this initialization complicates the quantization process. For instance, the all-zero distribution requires a separately designed quantizer at the beginning of tuning, since the scaling factor $s=0$ leads to an infinite $\frac{\mathbf{X}}{s}$ in Eq. 1. Second, the vanilla LoRA merges the FP16 $\mathbf{AB}$ and $\mathbf{W}$ using addition. When both $\mathbf{AB}$ and $\mathbf{W}$ are quantized, it is essential to ensure that they share identical quantization parameters to enable PTQ-free weight merging. This requirement leads to constrained parameter space, thus limiting the adaptation ability.